OpenStudy (mathwizzard3):

How is mass and gravity related to the speed of a falling object and the weight of that object?

OpenStudy (michele_laino):

hint: the lagrangian function \(\mathcal{L}\) of an object whose mass is \(m\), inside a gravitational field whose acceleration has magnitude \(g\), is: \[\huge \mathcal{L} = \frac{{m{{\dot z}^2}}}{2} - mgz\] now, please use such lagrangian function in order to write the equations of Lagrange: \[\huge - \frac{d}{{dt}}\frac{{\partial \mathcal{L}}}{{\partial \dot z}} + \frac{{\partial \mathcal{L}}}{{\partial z}} = 0\] furthermore, please keep in mind that we have: \[\huge v = \dot z\] where \(v\) is the magnitude of the speed of the object

OpenStudy (irishboy123):

and you can try this out!!! https://www.youtube.com/watch?v=5C5_dOEyAfk [m cancels out in the Lagrangian but the g doesn't.]

OpenStudy (anonymous):

I like maths!

OpenStudy (mathwizzard3):

Thank you everyone!